OpenStudy (anonymous):

Solve 7W= -21. A. 147 B. -147 C. 3 D. -3

OpenStudy (ajprincess):

divide both sides by 7

OpenStudy (anonymous):

divide both sides by 7 -21/7 = W

OpenStudy (ash2326):

We have \[ 7 w=-21\] Divide both sides by 7 \[\frac {7w}7=-\frac {21}{7}\] Can you solve now?

OpenStudy (anonymous):

You divied by -21/7 and you got -3